# Slant sums of quiver gauge theories

### v1

@@ 6.2 Proofs for ADE quivers
It follows from [29, Theorem 1.2] combined with [6, Remark 3.19] that the regular locus of $\overline{\mathcal{W}}^{\lambda}_{\mu}$ is equal to $\mathcal{W}^{\lambda}_{\mu}$. It remains to note that $z^{\mu}\in\mathcal{W}^{\lambda}_{\mu}$ iff $\mu\in W\lambda$. ∎
@@
Assume $\mu\in W\lambda$. For $\theta>0$, $\widetilde{\mathcal{M}}_{H}$ corresponds to the $\mu$ weight space of the irreducible representation with highest weight $\lambda$. Hence it is nonempty. Since $\dim\widetilde{\mathcal{M}}_{H}=(\lambda,\lambda)-(\mu,\mu)=0$, $\widetilde{\mathcal{M}}_{H}$ is a point. It is known, see for example Section 2.1 of [3], that the affinization of $\widetilde{\mathcal{M}}_{H}$ is independent of generic $\theta$. Hence $\widetilde{\mathcal{M}}_{H}$ is a point for any $\theta$.
@@
One can show that for an arbitrary quiver $Q$ without loops the corresponding quiver variety $\widetilde{\mathcal{M}}_{H}$ is a point iff $\mu\in W\lambda$. The argument is standard. Let us provide a sketch. The representation $V(\lambda)$ is integrable. Now, to every pair $\eta$, $\alpha_{i}$ of a weight $\eta$ of $V(\lambda)$ and a simple root $\alpha_{i}$ we consider the set of weights $S_{\eta,\alpha_{i}}$ of weights of $V(\lambda)$ of the form $\eta+k\alpha_{i}$ for $k\in\mathbb{Z}$. It is a standard fact (see [21, Proposition 3.6(b)]) that $\exists p,q,\in\mathbb{Z}_{\geq 0}$ such that $S_{\eta,\alpha_{i}}=\{\eta+k\alpha_{i}\,\mid\,-p\leq k\leq q\}$ and $s_{\alpha_{i}}(\eta-p\alpha_{i})=\eta+q\alpha_{i}$. Note now that $(\eta+k\alpha_{i},\eta+k\alpha_{i})=2k^{2}+2k(\eta,\alpha_{i})+(\eta,\eta)$ is a strictly convex function of $k$. So it must attain its maxima exactly at p-p and at $q$. Using this observation, one can prove that if $\mu$ is a weight of $V(\lambda)$, then $(\mu,\mu)\leqslant(\lambda,\lambda)$ and the equality holds iff $\mu\in W\lambda$ (argue by the induction on $\langle\rho,\mu\rangle$, where $\rho$ is the sum of fundamental weights for $\mathfrak{g}_{Q}$).
Proposition $6.$
Conjectures 5.2 and 5.3 hold for ADE quivers.
@@
Conjecture 5.2 follows from Lemmas 6.3, and 6.4. To prove Conjecture 5.3, recall that (18) identifies the partial resolution $\widetilde{\mathcal{M}}_{C,\nu}$ with the convolution diagram $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$. Now, Conjecture 5.3 follows from [25, Proposition 5.7] which states that the multiplication morphism $\tilde{\bf{m}}^{\underline{\lambda}}_{\underline{\mu}}$ is a $\mathsf{A}_{Q}$-equivariant open embedding. ∎
@@
Conjecture 1.5 holds for ADE quivers.
@@
The Coulomb branches we must consider are $\mathcal{M}_{C}=\overline{\mathcal{W}}^{\lambda}_{\mu}$ for $\mu\in W\lambda$. Recall the point $z^{\mu}$ lies in $\mathcal{W}^{\lambda}_{\mu}\subset\overline{\mathcal{W}}^{\lambda}_{\mu}$. Consider the repellent $R^{\lambda}_{\mu}$ (see [19, Definition 1.8.3]) with respect to the cocharacter $2\rho^{\vee}\colon\mathbb{C}^{\times}\rightarrow\mathsf{A}_{Q}$, given by the sum of the positive coroots, inside the smooth locus $\mathcal{W}^{\lambda}_{\mu}$. At the level of $\mathbb{C}$-points we have:
@@
Note that $R^{\lambda}_{\mu}$ is nonsingular (see [19, Proposition 1.7.6], or, alternatively, use Luna’s slice theorem) and $T_{z^{\mu}}R^{\lambda}_{\mu}=(T_{z^{\mu}}\overline{\mathcal{W}}^{\lambda}_{\mu})^{-}$, where - corresponds to taking the direct sum in $T_{z^{\mu}}\mathcal{W}^{\lambda}_{\mu}$ of negative $2\rho^{\vee}(\mathbb{C}^{\times})$-weight subspaces. Using the symplectic form on $\mathcal{W}^{\lambda}_{\mu}$ we obtain the isomorphism of representations of ${\mathsf{T}}_{Q}$:
@@
It remains to describe the ${\mathsf{T}}_{Q}$-character of $T_{z^{\mu}}R^{\lambda}_{\mu}$. It follows from [24, Theorem 3.1(1)] that $R^{\lambda}_{\mu}$ is isomorphic to the repellent to $z^{\mu}\in\operatorname{Gr}^{\lambda}$.
So, $T_{z^{\mu}}R^{\lambda}_{\mu}$ is isomorphic to $(T_{z^{\mu}}\operatorname{Gr}^{\lambda})^{-}$. We have
@@
and the claim follows from (20).
@@
Note that for $\lambda$ being minuscule and $\mu\in W\lambda$ Proposition 6.7 already follows from [25, Proposition 4.19], so the main new point of Proposition 6.7 is that the same formula holds for arbitrary dominant $\lambda$ and $\mu\in W\lambda$. It is worth noting that in [25, Section 4] a much stronger statement is proven for $\lambda$ minuscule and $\mu\in W\lambda$. Namely, it is proven that $\overline{\mathcal{W}}^{\lambda}_{\mu}=\mathcal{W}^{\lambda}_{\mu}\simeq T^{*}\mathbb{A}^{n}$ as symplectic varieties and the action of ${\mathsf{T}}_{Q}$ is described explicitly on the right hand side of the isomorphism. The same argument works for arbitrary $\lambda$ and $\mu\in W\lambda$ being almost dominant (i.e., $\langle\alpha,\mu\rangle\geqslant-1$ for any positive root $\alpha$).
Combining Propositions 6.6 and 6.7 we get the formula for the tangent space to an arbitrary nonsingular $\mathsf{T}_{Q}$-fixed point $z^{\underline{\mu}}$ of a partial resolution $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$ (recall that these points are in bijection with all possible decompositions $\mu=\mu_{1}+\ldots+\mu_{N}$ such that $\mu_{i}\in W\lambda_{i}$). This is a generalization of [25, Equation (5.8)]:
$T_{z^{\underline{\mu}}}\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}=\sum_{k=1}^{N}\sum_{\alpha\in\Phi^{-}_{\mu_{k}}}\sum_{i=1}^{\langle\alpha,\mu_{k}\rangle}\left(\hbar^{1-i-\langle\alpha,\mu_{1}+\ldots+\mu_{k-1}\rangle}e^{\alpha}+\hbar^{i+\langle\alpha,\mu_{1}+\ldots+\mu_{k-1}\rangle}e^{-\alpha}\right).$
@@ 6.3 About the restriction to real roots
Let us finish this section with a few remarks. When $Q$ is of type ADE, it is known that for $\mu$ being dominant $\overline{\mathcal{W}}^{\lambda}_{\mu}$ is indeed a slice to the $G_{\mathcal{O}}$-orbit $\operatorname{Gr}^{\mu}$ of $z^{\mu}$ inside the closure of the $G_{\mathcal{O}}$-orbit of $z^{\lambda}$. This is not the case for arbitrary $\mu$ (that is the reason why $\overline{\mathcal{W}}^{\lambda}_{\mu}$ are called generalized slices), although by [24, Theorem 3.1(1)] it is always true that the repellent to $z^{\mu}\in\overline{\mathcal{W}}^{\lambda}_{\mu}$ under the $\mathbb{C}^{\times}$-action coming from $2\rho^{\vee}$ coincides with the repellent to $z^{\mu}\in\overline{\operatorname{Gr}}^{\lambda}$.
Now assume that we are in the setting of Conjecture 1.5 (i.e., variety $\widetilde{\mathcal{M}}_{H}$ is a point), and we now allow $Q$ to be an arbitrary quiver without loops. This is equivalent to $\mu\in W\lambda$ (see Remark 6.5 above).
Assuming that the $2\rho^{\vee}$-fixed point $p\in\mathcal{M}_{C}$ is nonsingular and denoting by $R^{\lambda}_{\mu}\subset\mathcal{M}_{C}$ the $2\rho^{\vee}$-repellent to $p$ we see that Conjecture 1.5 is equivalent to the isomorphism of ${\mathsf{T}}_{Q}$-representations:
@@
where $\Phi^{-,\text{re}}$ are all real negative roots for $\mathfrak{g}=\mathfrak{g}_{Q}$. We do not know if the right hand side of (22) has a “geometric” meaning similar to the one in finite dimensional situation (when it is the tangent space to the repellent to $z^{\mu}\in\operatorname{Gr}_{G}^{\lambda}$). It is also not clear to us why only real roots appear in (22).
@@
Recall that $\mu\in W\lambda$ and pick $w\in W$ of minimal length such that $w(\lambda)=\mu$. It is a standard fact that $w$
has the following property: if $\beta$ is any positive root such that $\langle\beta,\lambda\rangle=0$, then $w\beta$ is also positive.
@@ 7 Slant sums of Coulomb branches
In this section, we study the slant sum construction from the Coulomb branch perspective.
@@ 7.1 Integrable systems for Coulomb branches
We have a natural (left) action of the algebra $H^{*}_{G_{\mathcal{O}}}(\operatorname{pt})=H^{*}_{G}(\operatorname{pt})$ on $\mathbb{C}[\mathcal{M}_{C}]=H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ as well as the action of $H^{*}_{G_{\mathcal{O}}\rtimes\mathbb{C}^{\times}}(\operatorname{pt})=H^{*}_{G\times\mathbb{C}^{\times}}(\operatorname{pt})$ on $\mathcal{A}$. Acting on identity, we obtain homomorphisms:
@@
defining integrable systems on $\mathcal{M}_{C}$ and $\mathcal{A}_{\hbar}$. In this way, we obtain (Poisson) commutative subalgebras of $\mathbb{C}[\mathcal{M}_{C}]$ and $\mathcal{A}$. The subalgebra $H:=H^{*}_{G\times\mathbb{C}^{\times}}(\operatorname{pt})|_{\hbar=1}\subset\mathcal{A}$ is called the Cartan subalgebra in [8]. It is also sometimes called Gelfand–Tsetlin subalgebra (see, for example, [42]). We can naturally identify it with the polynomial ring $\mathbb{C}[c_{i,k}\,|\,i\in Q_{0},\,k=1,\ldots,v_{i}]$, where $c_{i,k}=c_{i}^{G\times\mathbb{C}^{\times}}(V_{k})$ ($V_{k}$ is considered as a $G\times\mathbb{C}^{\times}$-equivariant bundle on $\operatorname{pt}$).
One reason why the algebra $H$ is important is because it allows one to define the notion of a Gelfand–Tsetlin character of a module over $\mathcal{A}$. Namely, whenever $M$ is a finitely generated $\mathcal{A}$-module such that the action of $H$ on it decomposes $M$ into the direct sum of finite dimensional generalized eigenspaces (such modules are called Gelfand–Tsetlin-modules in [42]), we can define its Gelfand–Tsetlin character as follows.
@@
where we treat $z^{\bf{S}}$ as formal parameters. For an ADE quiver $Q$, the Gelfand-Tsetlin character is one way to package the Frenkel-Reshetikhin $q$-character of a module. For short, we will write GT-character whenever we refer to the Gelfand-Tsetlin character.
7.2 Slant sum of Coulomb branches with one dimensional framing
Let $Q^{(1)}$ and $Q^{(2)}$ be quivers and fix dimension and framing vectors $\mathsf{v}^{(r)},\mathsf{w}^{(r)}$ for $r\in\{1,2\}$. We now make the following assumption.

### v2

@@ 7.2 Proofs for finite type Dynkin quivers
It follows from [undefal, Theorem 1.2] combined with [undefh, Remark 3.19] that the regular locus of $\overline{\mathcal{W}}^{\lambda}_{\mu}$ is equal to $\mathcal{W}^{\lambda}_{\mu}$. It remains to note that $z^{\mu}\in\mathcal{W}^{\lambda}_{\mu}$ iff $\mu\in W\lambda$. ∎
@@
Assume $\mu\in W\lambda$. For $\theta>0$, $\widetilde{\mathcal{M}}_{H}$ corresponds to the $\mu$ weight space of the irreducible representation with highest weight $\lambda$. Hence it is nonempty. Since $\dim\widetilde{\mathcal{M}}_{H}=(\lambda,\lambda)-(\mu,\mu)=0$, $\widetilde{\mathcal{M}}_{H}$ is a point. It is known, see for example Section 2.1 of [undefd], that the affinization of $\widetilde{\mathcal{M}}_{H}$ is independent of generic $\theta$. Hence $\widetilde{\mathcal{M}}_{H}$ is a point for any $\theta$.
@@
One can show that for an arbitrary quiver $Q$ without loops the corresponding quiver variety $\widetilde{\mathcal{M}}_{H}$ is a point iff $\mu\in W\lambda$. The argument is standard. Let us provide a sketch. The representation $V(\lambda)$ is integrable. Now, to every pair $\eta$, $\alpha_{i}$ of a weight $\eta$ of $V(\lambda)$ and a simple root $\alpha_{i}$ we consider the set of weights $S_{\eta,\alpha_{i}}$ of weights of $V(\lambda)$ of the form $\eta+k\alpha_{i}$ for $k\in\mathbb{Z}$. It is a standard fact (see [undefaa, Proposition 3.6(b)]) that $\exists p,q,\in\mathbb{Z}_{\geqslant 0}$ such that $S_{\eta,\alpha_{i}}=\{\eta+k\alpha_{i}\,\mid\,-p\leqslant k\leqslant q\}$ and $s_{\alpha_{i}}(\eta-p\alpha_{i})=\eta+q\alpha_{i}$. Note now that $(\eta+k\alpha_{i},\eta+k\alpha_{i})=2k^{2}+2k(\eta,\alpha_{i})+(\eta,\eta)$ is a strictly convex function of $k$. So it must attain its maxima exactly at p-p and at $q$. Using this observation, one can prove that if $\mu$ is a weight of $V(\lambda)$, then $(\mu,\mu)\leqslant(\lambda,\lambda)$ and the equality holds iff $\mu\in W\lambda$ (argue by the induction on $\langle\rho,\mu\rangle$, where $\rho$ is the sum of fundamental weights for $\mathfrak{g}_{Q}$).
Proposition 7.6.
Conjectures 5.1 and 5.3 hold for finite type Dynkin quivers.
@@
Conjecture 5.1 follows from Lemmas 7.3, and 7.4. To prove Conjecture 5.3, recall that (26) identifies the partial resolution $\widetilde{\mathcal{M}}_{C,\nu}$ with the convolution diagram $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$. Now, Conjecture 5.3 follows from [undefai, Proposition 5.7] which states that the multiplication morphism $\tilde{\bf{m}}^{\underline{\lambda}}_{\underline{\mu}}$ is a $\mathsf{A}_{Q}$-equivariant open embedding. ∎
@@
Conjecture 1.7 holds for finite type Dynkin quivers.
@@
The Coulomb branches we must consider are $\mathcal{M}_{C}=\overline{\mathcal{W}}^{\lambda}_{\mu}$ for $\mu\in W\lambda$. Recall the point $z^{\mu}$ lies in $\mathcal{W}^{\lambda}_{\mu}\subset\overline{\mathcal{W}}^{\lambda}_{\mu}$. Consider the repellent $R^{\lambda}_{\mu}$ (see [undefu, Definition 1.8.3]) with respect to the cocharacter $2\rho^{\vee}\colon\mathbb{C}^{\times}\rightarrow\mathsf{A}_{Q}$, given by the sum of the positive coroots, inside the smooth locus $\mathcal{W}^{\lambda}_{\mu}$. At the level of $\mathbb{C}$-points we have:
@@
Note that $R^{\lambda}_{\mu}$ is nonsingular (see [undefu, Proposition 1.7.6], or, alternatively, use Luna’s slice theorem) and $T_{z^{\mu}}R^{\lambda}_{\mu}=(T_{z^{\mu}}\overline{\mathcal{W}}^{\lambda}_{\mu})^{-}$, where - corresponds to taking the direct sum in $T_{z^{\mu}}\mathcal{W}^{\lambda}_{\mu}$ of negative $2\rho^{\vee}(\mathbb{C}^{\times})$-weight subspaces. Using the symplectic form on $\mathcal{W}^{\lambda}_{\mu}$ we obtain the isomorphism of representations of ${\mathsf{T}}_{Q}$:
@@
It remains to describe the ${\mathsf{T}}_{Q}$-character of $T_{z^{\mu}}R^{\lambda}_{\mu}$. It follows from [undefah, Theorem 3.1(1)] that $R^{\lambda}_{\mu}$ is isomorphic to the repellent to $z^{\mu}\in\operatorname{Gr}^{\lambda}$. So, $T_{z^{\mu}}R^{\lambda}_{\mu}$ is isomorphic to $(T_{z^{\mu}}\operatorname{Gr}^{\lambda})^{-}$. We have
@@
and the claim follows from (28). ∎
@@
Note that for $\lambda$ being minuscule and $\mu\in W\lambda$ Proposition 7.7 already follows from [undefai, Proposition 4.19], so the main new point of Proposition 7.7 is that the same formula holds for arbitrary dominant $\lambda$ and $\mu\in W\lambda$.
Combining Propositions 7.6 and 7.7 we get the formula for the tangent space to an arbitrary nonsingular $\mathsf{T}_{Q}$-fixed point $z^{\underline{\mu}}$ of a partial resolution $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$ (recall that these points are in bijection with all possible decompositions $\mu=\mu_{1}+\ldots+\mu_{N}$ such that $\mu_{i}\in W\lambda_{i}$). This is a generalization of [undefai, Equation (5.8)]:
$T_{z^{\underline{\mu}}}\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}=\sum_{k=1}^{N}\sum_{\alpha\in\Phi^{-}_{\mu_{k}}}\sum_{i=1}^{\langle\alpha,\mu_{k}\rangle}\left(\hbar^{1-i-\langle\alpha,\mu_{1}+\ldots+\mu_{k-1}\rangle}e^{\alpha}+\hbar^{i+\langle\alpha,\mu_{1}+\ldots+\mu_{k-1}\rangle}e^{-\alpha}\right).$ (30)
Note that for $\mu$ being dominant and without $\mathbb{C}^{\times}_{\hbar}$-equivariance, the equality (30) follows from [undefaf, Lemma 4.4]. We are grateful to Joel Kamnitzer for pointing this out to us.
@@ 7.3 About the restriction to real roots
Let us finish this section with a few remarks. When $Q$ is of finite type, it is known that for $\mu$ being dominant $\overline{\mathcal{W}}^{\lambda}_{\mu}$ is indeed a slice to the $G_{\mathcal{O}}$-orbit $\operatorname{Gr}^{\mu}$ of $z^{\mu}$ inside the closure of the $G_{\mathcal{O}}$-orbit of $z^{\lambda}$. This is not the case for arbitrary $\mu$ (that is the reason why $\overline{\mathcal{W}}^{\lambda}_{\mu}$ are called generalized slices), although by [undefah, Theorem 3.1(1)] it is always true that the repellent to $z^{\mu}\in\overline{\mathcal{W}}^{\lambda}_{\mu}$ under the $\mathbb{C}^{\times}$-action coming from $2\rho^{\vee}$ coincides with the repellent to $z^{\mu}\in\overline{\operatorname{Gr}}^{\lambda}$.
Now assume that we are in the setting of Conjecture 1.7 (i.e., variety $\widetilde{\mathcal{M}}_{H}$ is a point), and we now allow $Q$ to be an arbitrary quiver without edge loops. This is equivalent to $\mu\in W\lambda$ (see Remark 7.5 above). Assuming that the $2\rho^{\vee}$-fixed point $p\in\mathcal{M}_{C}$ is nonsingular and denoting by $R^{\lambda}_{\mu}\subset\mathcal{M}_{C}$ the $2\rho^{\vee}$-repellent to $p$ we see that Conjecture 1.7 is equivalent to the isomorphism of ${\mathsf{T}}_{Q}$-representations:
@@
where $\Phi^{-,\text{re}}$ are all real negative roots for $\mathfrak{g}=\mathfrak{g}_{Q}$. We do not know if the right hand side of (31) has a “geometric” meaning similar to the one in finite dimensional situation (when it is the tangent space to the repellent to $z^{\mu}\in\operatorname{Gr}_{G}^{\lambda}$).
Remark 7.9.
Recently, the paper [undeff] introduced schemes $\operatorname{Gr}^{\lambda}\cap T_{\mu}$ for arbitrary symmetrizable Kac-Moody group (see [undeff, Section 4.3]). It’s natural to expect that $R^{\lambda}_{\mu}$ is isomorphic to $\operatorname{Gr}^{\lambda}\cap T_{\mu}$ (for $Q$ of finite type this is known by [undefah, Theorem 3.1(1)]). One approach would be to extend the isomorphism $\mathcal{M}_{C}^{\mathrm{reg}}\simeq\mathcal{W}^{\lambda}_{\mu}$ of [undefh, Theorem 3.10] to this more general setting, where $\mathcal{W}^{\lambda}_{\mu}$ is constructed using [undeff, Section 4]. Once this identification is established, the same argument as in [undefah, Section 4.10] should yield the isomorphism $R^{\lambda}_{\mu}\simeq\operatorname{Gr}^{\lambda}\cap T_{\mu}$. Moreover, the isomorphism $\mathcal{M}_{C}^{\mathrm{reg}}\simeq\mathcal{W}^{\lambda}_{\mu}$ would imply Conjecture 1.7 in this greater generality (generalizing the finite type argument above) and hence verify the hypotheses used in Proposition 6.11. At present these ideas are completely speculative.
@@
Recall that $\mu\in W\lambda$ and pick $w\in W$ of minimal length such that $w(\lambda)=\mu$. It is a standard fact that $w$ has the following property: if $\beta$ is any positive root such that $\langle\beta,\lambda\rangle=0$, then $w\beta$ is also positive.
@@ 8 Slant sums of Coulomb branches
In this section, we begin the study of the slant sum construction from the Coulomb branch perspective. We focus on the case corresponding to Proctor’s original construction, namely $\mathsf{w}_{\star_{2}}^{(2)}=1$ and $\mathsf{w}_{i}^{(2)}=0$ for all other $i$. It would be very interesting (and much more useful for applications) to study the case of general $\mathsf{w}$ (see Section 8.2.2 above for some informal discussion).
@@ 8.1 Integrable systems for Coulomb branches
We have a natural (left) action of the algebra $H^{*}_{G_{\mathcal{O}}}(\operatorname{pt})=H^{*}_{G}(\operatorname{pt})$ on $\mathbb{C}[\mathcal{M}_{C}]=H_{*}^{G_{\mathcal{O}}}(\mathcal{R})$ as well as the action of $H^{*}_{G_{\mathcal{O}}\rtimes\mathbb{C}^{\times}}(\operatorname{pt})=H^{*}_{G\times\mathbb{C}^{\times}}(\operatorname{pt})$ on $\mathcal{A}_{\hbar}$. Acting on identity, we obtain homomorphisms:
@@
defining integrable systems on $\mathcal{M}_{C}$ and $\mathcal{A}_{\hbar}$. In this way, we obtain (Poisson) commutative subalgebras of $\mathbb{C}[\mathcal{M}_{C}]$ and $\mathcal{A}$. The subalgebra $H:=H^{*}_{G\times\mathbb{C}^{\times}}(\operatorname{pt})|_{\hbar=1}\subset\mathcal{A}$ is called the Cartan subalgebra in [undefj]. It is also sometimes called Gelfand–Tsetlin subalgebra (see, for example, [undefaz]). We can naturally identify it with the polynomial ring $\mathbb{C}[c_{i,k}\,|\,i\in Q_{0},\,k=1,\ldots,v_{i}]$, where $c_{i,k}=c_{i}^{G\times\mathbb{C}^{\times}}(V_{k})$ ($V_{k}$ is considered as a $G\times\mathbb{C}^{\times}$-equivariant bundle on $\operatorname{pt}$).
One reason why the algebra $H$ is important is because it allows one to define the notion of a Gelfand–Tsetlin character of a module over $\mathcal{A}$. Namely, whenever $M$ is a finitely generated $\mathcal{A}$-module such that the action of $H$ on it decomposes $M$ into the direct sum of finite dimensional generalized eigenspaces (such modules are called Gelfand–Tsetlin-modules in [undefaz]), we can define its Gelfand–Tsetlin character as follows.
@@
where we treat $z^{\bf{S}}$ as formal parameters. For a finite type Dynkin quiver $Q$, the Gelfand-Tsetlin character is one way to package the Frenkel-Reshetikhin $q$-character of a module.
8.2 Slant sum of Coulomb branches
Let $Q^{(1)}$ and $Q^{(2)}$ be quivers and fix dimension and framing vectors $\mathsf{v}^{(r)},\mathsf{w}^{(r)}$ for $r\in\{1,2\}$.
8.2.1 One-dimensional framing
We now make the following assumption.